# From Fragments to Facts: A Curriculum-Driven DPO Approach for Generating Hindi News Veracity Explanations

### v2

@@ Abstract
In an era of rampant misinformation, generating reliable news explanations is vital, especially for under-represented languages like Hindi. Lacking robust automated tools, Hindi faces challenges in scaling misinformation detection. To bridge this gap, we propose a novel framework integrating Direct Preference Optimization (DPO) with curriculum learning to align machine-generated explanations with human reasoning. Fact-checked explanations from credible sources serve as preferred responses, while LLM outputs highlight system limitations and serve as non-preferred responses. To refine task-specific alignment, we introduce two key parameters—Actuality and Finesse—into the DPO loss function, enhancing explanation quality and consistency. Experiments with LLMs (Mistral, Llama, Gemma) and PLMs (mBART, mT5) confirm the framework’s effectiveness in generating coherent, contextually relevant explanations. This scalable approach combats misinformation and extends automated explanation generation to low-resource languages.
@@ 1 Introduction
The rise of fake news, fuelled by its low production cost and widespread digital dissemination, poses a significant threat to the integrity of journalism. As Toomas Hendrik Ilves aptly states, “Fake news is cheap to produce. Genuine journalism is expensive." This disparity is evident in the societal disruptions caused by misinformation, particularly during crises like the COVID-19 pandemic, where unverified claims about treatments and preventive measures fuelled panic and confusion (Barua et al., 2020; Roozenbeek et al., 2020). Misinformation also exacerbates political polarization, creating deep societal divides (Cantarella et al., 2023; Bovet & Makse, 2019). Fact-checking platforms, while essential, face substantial challenges in scaling their efforts, especially in languages like Hindi. Despite over 600 million Hindi speakers [^1] https://www.britannica.com/topic/languages-by-total-number-of-speakers-2228881, automated tools for generating credible, human-like explanations in Hindi remain underdeveloped. Addressing this gap is crucial to supporting fact-checking initiatives and reducing the impact of fake news on society.
Overview: To address these challenges, this paper introduces the DeFactoX framework, which evaluates the veracity of Hindi news articles—while remaining adaptable to other languages—and generates coherent, factually grounded explanations supporting its prediction. By analyzing the content, DeFactoX determines whether a news piece is credible or misleading and provides a well-reasoned justification, enhancing transparency and trust in automated misinformation detection.
We create a synthetic Hindi preference dataset by using fact-checked explanations written by humans as preferred responses and machine-generated outputs from various LLMs as rejected responses. To enhance the generation of high-quality explanations, we improve the Direct Preference Optimization (DPO) approach (Rafailov et al., 2024) by integrating fact scores, variance-based reduction parameters, and curriculum learning. These enhancements align generated explanations with human reasoning, ensuring reliability and accuracy. Leveraging curriculum learning’s data augmentation, our scalable framework adapts to low-resource languages, aiding global misinformation detection and automated explanation generation.
@@
Pre-trained LLMs, trained on generalized datasets, struggle to assess the veracity of Hindi news and generate contextually relevant, factually grounded explanations for the veracity predicted. Moreover, fact-checking in Hindi remains predominantly manual, lacking scalable automated solutions. Given Hindi’s vast speaker base and the increasing spread of misinformation, it is crucial to develop robust, scalable methods for automated veracity prediction and explanation generation. While our framework is language-agnostic, we focus on Hindi to address this pressing need and demonstrate its effectiveness in a low-resource setting.
@@
  • RQ-1: How can automated systems reliably assess the veracity of Hindi news and generate human-like explanations that are coherent, contextually relevant, and factually accurate while explicitly justifying the model’s veracity predictions?
  • RQ-2: Can Direct Preference Optimization (DPO) (Rafailov et al., 2024), augmented with FactScore (Min et al., 2023) and variance-based parameters, effectively align machine-generated explanations with human standards in resource-constrained languages like Hindi, ensuring both factual accuracy and interpretability?
  • RQ-3: How can curriculum learning (Pattnaik et al., 2024) be integrated with DPO to refine veracity prediction in under-represented languages, and what scalable methodologies can extend misinformation detection and explanation generation to other low-resource languages?
Research Motivation: The growing spread of misinformation in underrepresented languages like Hindi underscores the urgent need for scalable, automated systems that can assess veracity and generate reliable, human-like explanations. Unlike high-resource languages, Hindi suffers from a lack of robust fact-checking tools, and existing LLMs often fall short of maintaining coherence, factual accuracy, and human alignment for veracity explanations.
This research aims to bridge these gaps by leveraging Direct Preference Optimization (DPO) (Rafailov et al., 2024), curriculum learning (Pattnaik et al., 2024), FactScore (Min et al., 2023), and variance-based reduction techniques to refine veracity explanation generation, ensuring both accuracy and scalability in combating misinformation.
Contributions The key contributions of our research are as follows:
  • We introduce a synthetic, ranking-based Hindi preference dataset, where human-written fact-checked veracity explanations serve as the highest-ranked responses, while machine-generated outputs from various LLMs are evaluated and ranked based on a weighted scoring mechanism incorporating BERTSCORE (Zhang et al., 2019), ROUGE-L (Lin, 2004), and METEOR (Banerjee & Lavie, 2005). This ensures that while machine-generated explanations are treated as non-preferred responses, they are ranked based on quality rather than being uniformly categorized as poor.
  • We introduce an enhanced DPO framework, DeFactoX, inspired by the principle of clear, fact-based, human-written veracity explanations. Our framework integrates FactScore, variance-based reduction parameters, and curriculum learning to train models that not only predict veracity but also generate explanations that are coherent, contextually accurate, and aligned with human reasoning, thereby strengthening the model’s veracity claims.
  • Our language-agnostic framework lays the groundwork for automated veracity-driven explanation generation in low-resource languages. Focusing on Hindi due to its underrepresentation in fact-checking, we present the first framework of its kind, contributing to the global effort to combat misinformation with a scalable and adaptable solution.
@@ 2 Related Works
We explore the growth of misinformation detection through veracity explanation alongside advancements in Preference Optimization techniques.
@@
Automated Misinformation Detection and Explanation Generation: Recent studies have advanced misinformation detection and explanation generation. Joshi et al. (2023) integrated Domain Adversarial Neural Networks (DANN) with LIME to enhance COVID-19 misinformation detection. Chi & Liao (2022) proposed QA-AXDS, a scalable, interpretable fake news detection system using dialogue trees. Yao et al. (2023) introduced MOCHEG, a multimodal fact-checking benchmark incorporating textual and visual evidence. Zhou et al. (2023) examined AI-generated misinformation, highlighting linguistic nuances and proposing updated detection guidelines. Gong et al. (2024) emphasized socio-contextual cues in “social explanation” to combat misinformation. Russo et al. (2023) showed that extractive steps improve abstractive summarization for claim verification. Bilal et al. (2024) developed a GNN-based rumour verification model leveraging opinion-guided summaries. Yue et al. (2024) introduced RARG, combining evidence retrieval with RLHF-tuned LLMs, excelling in COVID-19 misinformation detection. The study by Yang et al. (2022) employed a QA-based framework with attention-driven comparisons for interpretable fact-checking. For a comprehensive review, readers are pointed to the work by Kotonya & Toni (2020). Our Novelty: While previous works primarily target high-resource languages like English and Chinese, our focus is on Hindi, an under-represented language.
@@
Applications and Advancements in Preference Optimization and Curriculum Learning: Recent advancements in preference optimization and curriculum learning have enhanced model performance across domains. Pattnaik et al. (2024) introduced Curry-DPO, a curriculum learning-based enhancement of DPO, achieving up to 7.5% improvement across datasets. Chen et al. (2024a) proposed a multi-stage curriculum framework optimizing humour and structure preferences in LLMs. Yin et al. Yin et al. (2024) developed Self-Augmented Preference Optimization (SAPO), surpassing DPO and SPIN across multiple benchmarks. Morimura et al. (2024) introduced filtered DPO (fDPO), refining datasets for better training efficiency. Wang et al. (2024) proposed Balanced Preference Optimization (BPO), enhancing knowledge depth while maintaining efficiency. Zeng et al. (2024) presented Token-DPO, improving alignment and diversity in LLMs through token-level fine-tuning. Wallace et al. (2024) introduced Diffusion-DPO, aligning diffusion models with human preferences for superior visual appeal. Chen et al. (2024b) proposed Softmax-DPO to enhance recommender systems via user preference modelling. Lai et al. (2024) introduced Step-DPO, improving mathematical reasoning in LLMs with minimal data. For a comprehensive review, readers are referred to the study by Xiao et al. (2024). Our Novelty: While DPO has been applied across various domains, our approach is specifically tailored to generating veracity claims and explanations for Hindi news. The research gap in this area highlights challenges faced by under-represented languages like Hindi, including data scarcity and the limitations of LLMs trained on multilingual datasets. Our study offers a scalable and effective solution to address these challenges. Remark: A detailed section on the theoretical foundations of DPO (Rafailov et al., 2024) and Curry-DPO (Pattnaik et al., 2024) is provided in the appendix (section A.5 & A.6) for interested readers.
Figure 1: Overview of DeFactoX framework.
@@ 3.1 Dataset Selection and Sampling
Multiple veracity claim misinformation detection datasets Bhardwaj et al. (2020); Kumar & Singh (2022); Sharma & Arya (2024); Bansal et al. (2024); Sharma & Garg (2021); Badam et al. (2022); Sharma & Arya (2023) provide Hindi news articles sourced from fact-checking websites, ensuring authentic veracity labels. We selected data instances from Sharma & Arya (2024), a comprehensive dataset featuring over 15,000 articles in the fake news category and 13,000+ in the real news category. This dataset was chosen for its extensive coverage of news, spanning from older to recent events, and its sourcing from fact-checking websites, which provide verified veracity labels. To maintain a balanced and manageable dataset, we extracted the most recent 3,000 articles from each class (fake and real news). The selection ensures a healthy mix of data samples without being overbearing in size.
@@ 3.2 Characteristics of the Selected Data
(1) Each article is sourced from reputable fact-checking websites that not only classify news as fake or real but also provide comprehensive, well-reasoned explanations justifying these classifications. These explanations serve as ground truth references for evaluating model-generated outputs.
(2) The first paragraph of every article strictly contains only the core news content, intentionally excluding any veracity indicators or reasoning. This neutral presentation ensures that readers, and more importantly, models—cannot determine whether the news is fake or real based solely on this segment.
@@
3.3 Observations on Explanations
A key observation in our dataset is the distinct difference in the writing styles of explanations for true and fake news.
(1) True news explanations: Fact-checking sources often offer minimal explanations for true news, assuming that the facts and details presented are self-evident. These explanations are typically brief and rarely include explicit statements affirming the news’ authenticity.
(2) Fake news explanations: In contrast, explanations for fake news are far more detailed and explicit. Fact-checkers provide strong declarative statements, rejecting falsehoods and supplementing them with clear justifications, such as evidence-based counterarguments, source verification, and logical reasoning. This explicit mention of veracity is illustrated in Figure 2. To gain further insight into the nature of these explanations, readers are encouraged to refer to the original fact-checking sources, including OneIndia [^2] https://hindi.oneindia.com/fact-check/, Vishvas News [^3] https://www.vishvasnews.com/, and Aaj Tak [^4] https://www.aajtak.in/fact-check, which serve as the primary references for this dataset.
@@ 3.4 Standardizing True News Explanations
While explanations for fake news naturally include explicit reasoning and veracity statements, true news explanations often lack such clarity. Fact-checking sources assume that factual news is self-evident, leading to minimal justifications. This inconsistency makes it difficult for models to learn a uniform veracity-based explanation structure. To address this, we standardized true news explanations by ensuring they explicitly affirm their veracity while preserving factual integrity. This step aligns true news explanations with the structured reasoning seen in fake news explanations, thereby creating a more balanced training dataset. For implementation details and examples, including the standardization prompt, please refer to the appendix (Section A.2 & A.3).
@@ 3.5 Generating Rejected Responses
To construct the negative class (non-preferred or rejected responses), we used the first paragraph of each article, presenting only the core news content, whether fake or true, as input for three state-of-the-art LLMs: gpt-4o-mini Achiam et al. (2023), Mistral-7B-v0.1 Jiang et al. (2023), and gemini-1.5-flash Team et al. (2024). These models were selected due to their strong reasoning capabilities and proven performance in NLP tasks (Liu et al., 2024; Mathur et al., 2024; Siino, 2024a; b; Trott & Rivière, 2024; Sato et al., 2024).
Hallucination Mitigation: Numerous recent studies Rosset et al. (2023); Hu et al. (2023); Ghosh et al. (2024) have leveraged language models to generate synthetic preference datasets for reinforcement learning via human feedback (RLHF), inspiring our approach as well. A primary concern in generating synthetic explanations is the risk of hallucination—incorrect or fabricated details in the model’s output. To mitigate hallucination, we designed a controlled prompt that explicitly restricted responses to the provided news content, minimizing the inclusion of extraneous information.
Additionally, we manually reviewed 1000 randomly sampled explanations from both classes, verifying factual accuracy. Key checks included identifying geographical inconsistencies (e.g., misrepresenting an Indian event as occurring in the USA), verifying temporal accuracy (ensuring dates and timelines matched the original article), detecting entity distortions (confirming that names of people, organizations, and locations were correctly retained), and assessing logical coherence (ensuring explanations aligned with the core facts without introducing contradictions). No significant errors were found across these dimensions, reinforcing the overall reliability of our dataset. For implementation details, including the full prompt used for generating rejected responses, readers are pointed to the appendix (section  A.4).
We present the algorithmic overview of Dataset Creation from Fact-Checked News Articles in Algorithm 1

### v3

@@ Abstract
In an era of rampant misinformation, generating reliable news explanations is vital, especially for underrepresented languages like Hindi. Lacking robust automated tools, Hindi faces challenges in scaling misinformation detection. To bridge this gap, we propose DeFactoX, a novel framework integrating Direct Preference Optimization (DPO) with Curriculum learning to align machine-generated explanations with human reasoning. Fact-checked explanations from credible sources serve as preferred responses, while LLM outputs highlight system limitations and serve as non-preferred responses. At the core of this framework lies Hin-DPO, an enhanced variant of DPO that enriches the loss function with two novel parameters, Actuality and Finesse, enhancing explanation quality and consistency. Experiments with LLMs (Mistral, Llama, Gemma) and PLMs (mBART, mT5) confirm the framework’s effectiveness in generating coherent, contextually relevant explanations.
From Fragments to Facts: A Curriculum-Driven DPO Approach for Generating Hindi News Veracity Explanations
Pulkit Bansal1   Raghvendra Kumar1   Shakti Singh1   Sriparna Saha1   Adam Jatowt2 1Indian Institute of Technology Patna, India 2University of Innsbruck, Austria {pulkit_2101mc48, raghvendra_2221cs27}@iitp.ac.in
@@ 1 Introduction
The rise of fake news, fuelled by its low production cost and widespread digital dissemination, poses a significant threat to the integrity of journalism. As Toomas Hendrik Ilves aptly states, “Fake news is cheap to produce. Genuine journalism is expensive." This disparity is evident in the societal disruptions caused by misinformation, particularly during crises like the COVID-19 pandemic, where unverified claims about treatments and preventive measures fuelled panic and confusion (Barua et al., 2020; Roozenbeek et al., 2020). Misinformation also exacerbates political polarization, creating deep societal divides (Cantarella et al., 2023; Bovet and Makse, 2019). Fact-checking platforms face substantial challenges in scaling their efforts, especially in languages like Hindi. Despite over 600 million Hindi speakers [^1] https://www.britannica.com/topic/languages-by-total-number-of-speakers-2228881, automated tools for generating credible, human-like explanations in Hindi remain underdeveloped. Addressing this gap is crucial to supporting fact-checking initiatives and reducing the impact of fake news on society.
Overview: This paper introduces DeFactoX, a framework for Hindi news veracity prediction and explanation generation. The framework first constructs a synthetic preference dataset in which human-written, fact-checked explanations serve as preferred responses and LLM-generated explanations serve as rejected responses, thereby grounding alignment in reliable human reasoning while exposing the model to common erroneous patterns.
Building on this dataset, we use Curriculum learning (Pattnaik et al., 2024) to progressively train the model to distinguish between explanations of increasing difficulty. To further strengthen alignment, we extend Direct Preference Optimization (DPO) (Rafailov et al., 2024) with a novel loss function, termed Hin-DPO, which incorporates two additional parameters: Actuality, quantifying the factual correctness of responses, and Finesse, measuring hallucination through output instability across generations. Together, these innovations enable DeFactoX to generate explanations that are factually accurate, consistent, and trustworthy.
@@
Pre-trained LLMs, trained on generalized datasets, struggle to assess the veracity of Hindi news and generate contextually relevant, factually grounded explanations for the veracity predicted. Moreover, fact-checking in Hindi remains predominantly manual, lacking scalable automated solutions. Given Hindi’s vast speaker base and the increasing spread of misinformation, it is crucial to develop robust, scalable methods for automated veracity prediction and explanation generation.
@@
  • RQ-1: How can automated systems reliably assess the veracity of Hindi news and generate human-like explanations that are coherent and contextually relevant?
  • RQ-2: How effective is incorporating parameters such as Actuality (factual accuracy) and Finesse (hallucination sensitivity) in aligning model outputs with human preferences?
  • RQ-3: How can Curriculum Learning be combined with DPO to progressively enhance veracity prediction and explanation generation for Hindi news?
Research Motivation: The rapid spread of misinformation in languages like Hindi highlights the need for scalable systems that assess veracity and generate reliable, human-like explanations. Unlike high-resource languages, Hindi lacks robust fact-checking tools, and existing LLMs often struggle with coherence, factual accuracy, and human alignment, motivating the need for novel frameworks that ensure trustworthy explanations.
This work addresses these gaps by refining veracity explanation generation through Direct Preference Optimization (DPO) (Rafailov et al., 2024), Curriculum learning (Pattnaik et al., 2024), and our enhanced loss function Hin-DPO, which integrates the Actuality score (inspired by FactScore (Min et al., 2023)) and the Finesse score (a variance-based measure of hallucination), ensuring both accuracy and scalability.
Contributions:
  • We create a synthetic, ranking-based Hindi preference dataset, where human fact-checked explanations serve as top-ranked responses, and LLM outputs are ranked using an automated scoring mechanism that closely aligns with human preferences.
  • We propose DeFactoX, a two-stage framework that combines curriculum learning with an enhanced preference optimization objective, Hin-DPO, which integrates two novel parameters: Actuality and Finesse.
  • To the best of our knowledge, DeFactoX is the first framework for automated veracity-driven explanation generation in Hindi. Its data augmentation based design provides a scalable methodology that can facilitate future adaptation to other languages.
Figure 1: Overview of DeFactoX framework. (Left) Dataset creation with human-written explanations as preferred responses and LLM-generated explanations as rejected responses. (Center) Curriculum-based dataset construction, where samples are ranked and bucketed into easy, moderate, and hard levels. (Right) Model training with Hin-DPO under curriculum learning, fine-tuning the model to generate human aligned explanations.
@@ 2 Related Works
Automated Misinformation Detection and Explanation Generation: Recent studies have advanced misinformation detection and explanation generation. Joshi et al. (2023) integrated Domain Adversarial Neural Networks (DANN) with LIME to enhance COVID-19 misinformation detection. Chi and Liao (2022) proposed QA-AXDS, a scalable, interpretable fake news detection system using dialogue trees. Yao et al. (2023) introduced MOCHEG, a multimodal fact-checking benchmark incorporating textual and visual evidence. Zhou et al. (2023) examined AI-generated misinformation, highlighting linguistic nuances and proposing updated detection guidelines. Gong et al. (2024) emphasized socio-contextual cues in “social explanation” to combat misinformation. Russo et al. (2023) showed that extractive steps improve abstractive summarization for claim verification. Bilal et al. (2024) developed a GNN-based rumour verification model leveraging opinion-guided summaries. Yue et al. (2024) introduced RARG, combining evidence retrieval with RLHF-tuned LLMs, excelling in COVID-19 misinformation detection. The study by Yang et al. (2022) employed a QA-based framework with attention-driven comparisons for interpretable fact-checking. For a comprehensive review, readers are pointed to the work by Kotonya and Toni (2020).
Our Novelty: While previous works primarily target high-resource languages like English and Chinese, our focus is on Hindi, an under-represented language.
@@
Applications and Advancements in Preference Optimization and Curriculum Learning: Recent advancements in preference optimization and Curriculum learning have enhanced model performance across domains. Pattnaik et al. (2024) introduced Curry-DPO, a Curriculum learning-based enhancement of DPO, achieving up to 7.5% improvement across datasets. Chen et al. (2024a) proposed a multi-stage Curriculum framework optimizing humour and structure preferences in LLMs. Yin et al. (2024) developed Self-Augmented Preference Optimization (SAPO), surpassing DPO and SPIN across multiple benchmarks. Morimura et al. (2024) introduced filtered DPO (fDPO), refining datasets for better training efficiency. Wang et al. (2024) proposed Balanced Preference Optimization (BPO), enhancing knowledge depth while maintaining efficiency. Zeng et al. (2024) presented Token-DPO, improving alignment and diversity in LLMs through token-level fine-tuning. Chen et al. (2024b) proposed Softmax-DPO to enhance recommender systems via user preference modelling. Lai et al. (2024) introduced Step-DPO, improving mathematical reasoning in LLMs with minimal data. For a comprehensive review, readers are referred to the study by Xiao et al. (2024).
Our Novelty: While DPO has been applied across various domains, our approach is specifically tailored to generating veracity claims and explanations for Hindi news. The research gap in this area highlights challenges faced by Hindi, including data scarcity and the limitations of LLMs trained on multilingual datasets. Our study offers an effective solution to address these challenges.
@@ 3.1 Dataset Selection and Sampling
Multiple veracity claim misinformation detection datasets Bhardwaj et al. (2020); Kumar and Singh (2022); Sharma and Arya (2024); Bansal et al. (2024); Badam et al. (2022); Sharma and Arya (2023); Kumar et al. (2025) provide Hindi news articles sourced from fact-checking websites, ensuring authentic veracity labels. We selected data instances from Sharma and Arya (2024), a comprehensive dataset featuring over 15,000 articles in the fake news category and 13,000+ in the real news category. This dataset was chosen for its extensive coverage of news, spanning from older to recent events, and its sourcing from fact-checking websites, which provide verified veracity labels. To maintain a balanced and manageable dataset, we extracted the most recent 5,000 articles from each class (fake and real). The selection ensures a healthy mix of data samples without being overbearing in size. Further stats are presented in the appendix (Section A.8).
@@ 3.2 Characteristics of the Selected Data
(1) Each article is sourced from fact-checking websites that not only classify news as fake or real but also provide comprehensive, well-reasoned explanations justifying these classifications. These explanations serve as ground truth references for evaluating model-generated outputs.
(2) The first paragraph of every article strictly contains only the core news content, intentionally excluding any veracity or reasoning. This neutral presentation ensures that readers, and more importantly, models cannot determine whether the news is fake or real based solely on this segment.
@@
3.3 Observations on Explanations
A key observation in the dataset is the distinct difference in the writing styles of explanations for true and fake news.
@@ 3.3 Observations on Explanations
(1) True news explanations: During our manual verification, we observed that many human-written explanations were primarily informational in nature. They tended to summarize the facts of the news story but did not include explicit reasoning or assertive statements confirming its authenticity. Such explanations typically present the news content in a descriptive manner, without additional justification or emphasis on truthfulness. For an example, see appendix (Section A.2).
(2) Fake news explanations: In contrast, explanations for fake news are far more detailed and explicit. Fact-checkers provide strong declarative statements, rejecting falsehoods and supplementing them with clear justifications, such as evidence-based counterarguments, source verification, and logical reasoning. This explicit mention of veracity is illustrated in Figure 2. For a deeper understanding of these explanations, readers may consult the original fact-checking sources: OneIndia [^2] https://hindi.oneindia.com/fact-check/, Vishvas News [^3] https://www.vishvasnews.com/, and Aaj Tak [^4] https://www.aajtak.in/fact-check, which serve as the primary references for this dataset and are certified by the International Fact-Checking Network (IFCN).
@@ 3.4 Standardizing True News Explanations
While explanations for fake news naturally include explicit reasoning and veracity statements, true news explanations often lack such clarity. Fact-checking sources assume that factual news is self-evident, leading to minimal justifications. This inconsistency makes it difficult for models to learn a uniform veracity-based explanation structure. To address this, we standardized true news explanations by ensuring they explicitly affirm their veracity while preserving factual integrity. This step aligns true news explanations with the structured reasoning seen in fake news explanations. For implementation details and examples, please refer to the appendix (Section A.3 & A.6).
@@ 3.5 Generating Rejected Responses
To construct the rejected responses, we used the first paragraph of each article, presenting only the core news content, as input for three state-of-the-art LLMs: gpt-4o-mini Achiam et al. (2023), Mistral-7B-v0.1 Jiang et al. (2023), and gemini-1.5-flash Team et al. (2024). These models were selected due to their strong reasoning capabilities and proven performance in NLP tasks (Liu et al., 2024; Mathur et al., 2024; Siino, 2024a, b; Trott and Rivière, 2024; Sato et al., 2024).
Unlike existing approaches that aim to minimize hallucinations in synthetic preference datasets, our goal is fundamentally different: we intentionally generated weaker, non-preferred explanations using the prompt in appendix (Section A.4). We use a simple prompt without additional fine-tuning or safeguards, letting the generated responses naturally reflect the inherent limitations of LLMs.
Observations: Through manual verification of 1,500 sampled rejected responses, we found recurring issues that distinguished them from human-written explanations. Specifically, LLM-generated explanations often:
  • Over-emphasized superficial linguistic elements or unusual words in the news text.
  • Displayed bias towards stylistic or surface-level framing rather than factual grounding.
  • Focused on a single aspect of news instead of covering multiple factual dimensions.
  • Lacked the depth, balance, and context-awareness consistently present in human-authored explanations.
These weaknesses confirmed their suitability as the negative class for preference optimization.